OpenStudy (anonymous):

4x-8y=8 find the equation in point-slope form, of the line that is perpendicular to this line and passes through the point (1,2)

OpenStudy (anonymous):

sope of the line ax+by=c is given by -a/b hence the reqd slope here is -4/-8=1/2 the slope of the line perp. to the above will have slope of -(1/(1/2))=-2

OpenStudy (anonymous):

but th above line passes through thus the reqd line in point-slope form, is given by (y-2) = -2 (x-1)

OpenStudy (anonymous):

so its (-8y-2)=-2(4x-1) but what is the line perpendicular to that line and passes through(1,2)?